# Information-Driven Design of Imaging Systems

### v4

@@ Acknowledgements
We conducted experiments using the Gehler-Shi dataset [54, 55] (implied license via explicit permission to use), which comprises 568 high-quality natural images. The dataset was partitioned into 461 training, 51 validation, and 56 test images, with each image subdivided into $24\times 24$ pixel patches. To establish a consistent baseline for illumination, we computed a white channel by summing the red, green, and blue channels, then normalized the dataset to achieve a mean of $1000$ photons per pixel in the white channel, based on the training split mean. Signal-dependent Poisson noise was subsequently applied to all patches.
@@
For mutual information estimation, we sampled $100,000$ patches from the test set for fitting a PixelCNN, reserving $10,000$ for evaluation. Conditional entropy calculations were performed on clean images prior to synthetic Poisson noise application. We evaluated three distinct color filter array (CFA) designs: a Bayer pattern, a random configuration, and a learned pattern using the architecture in [53]. For each CFA design, we selected the minimum information estimate across the 20 replicates (since our estimator provides an upper bound) and the best prediction performance (lowest negative log likelihood on test data) to characterize system performance. Each replicate took $\sim$3.5GB of memory. All replicates were trained in $\sim$5 hours on an Nvidia RTX A6000 GPU.
@@
Color image reconstruction was implemented using a bifurcated network architecture as proposed by Chakrabarti [53]. This approach processes $24\times 24$ pixel measurement patches through two parallel paths. The first path performs multiplicative operations in log-space with linear mixing, while the second employs convolutional layers with $128$ filters. Each path generates $24$ estimates per color value, which are then combined through learned weights for final reconstruction.
@@
For our site selection experiments, we generated a dataset of 100,000 synthetic black holes based on the physical models for black hole dynamics from [147]. Each synthetic black hole had a static envelope, right ascension, and source declination matching that of M87. The dynamic portion forming the accretion disk surrounding each black hole was modeled as a Gaussian random field generated by a different solution to an underlying anisotropic spatio-temporal diffusion partial differential equation with fixed diffusion and advection fields. Each black hole was formed by applying a Gaussian random field as a multiplicative perturbation to the static envelope, and normalized to have total flux of 1 Jansky. To simulate black hole imaging we selected combinations of four telescopes from the eight original telescope locations used in the Event Horizon Telescope 2017 array. We also included comparisons between reconstruction quality and information estimates for all 70 possible combinations of four telescopes (Fig. S20). For each combination we simulated radio telescope measurements including thermal noise using the methods described in [148].
@@
Black hole images were reconstructed from simulated radio telescope measurements using a standard regularized maximum likelihood method with total variation regularization [148], with hyperparameters matching the implementation in [149]. Images were reconstructed for a 2,000 measurement subset of the test set used for information estimation. Each reconstruction was performed on the CPU and took $\sim 10$s.
@@
The CIFAR10 dataset [150] (implied license via permission to use) was used for the lensless imaging experiments. This dataset consists of 60,000 total images across 10 classes. To convert CIFAR10 image pixel values to synthetic photon counts, the dataset was scaled to have mean value equivalent to the desired photons per pixel. Encoded measurements for each lensless imaging system were generated by convolving (valid region, no zero padding) with the corresponding point spread function. The point spread functions for the one lens and four lens encoders were modeled as Gaussian-blurred points. For the diffuser encoder, the point spread function is a downsampled version of an experimentally captured point spread function from DiffuserCam [59]. After convolution, synthetic Poisson noise was added to form the noisy encoded measurements.
@@
We analyzed single leukocyte images and corresponding protein expression measurements from the Berkeley Single Cell Computational Microscopy (BSCCM) dataset [65] (CC0 1.0 Universal license). The dataset includes images acquired under multiple illumination conditions using an LED array microscope along with measurements of eight protein markers (CD3, CD19, CD56, CD123, HLA-DR, CD14, CD16, and CD45) obtained through antibody staining. To study performance at different signal levels, we simulated lower photon counts by adding synthetic Poisson noise to the original images.
@@
We developed a neural network to predict protein expression levels from single-cell images. The network uses a DenseNet121 [151] backbone pretrained on ImageNet, modified to accept single-channel input by averaging the first convolutional layer’s RGB weights. The backbone feeds into a global average pooling layer followed by eight parallel fully-connected networks, each with two hidden layers. These networks output parameters for Gaussian mixture models corresponding to each protein marker, enabling direct evaluation of the negative log likelihood for target protein levels. Training used the Adam optimizer with a learning rate of $5\times 10^{-5}$ and batches of 16 images. A composite loss function summed the negative log likelihood across protein markers, with missing values masked out. Training proceeded for 4,000 steps per epoch and employed early stopping when validation loss failed to improve for 20 epochs. Typical trainings took 1-2 days on NVIDIA TITAN Xp GPUs with 12GB memory. 2-3 networks were trained in parallel sharing a single GPU.
@@
S6.6 Information Driven Encoder Analysis Learning (IDEAL)
Using the Gehler-Shi dataset [54, 55] partitioned into training (461 images), validation (51 images), and test (56 images) sets, we extracted 100,000 patches of size $24\times 24$ pixels for training, along with 10,000 patches from the validation set. To establish consistent illumination conditions, the dataset was normalized to achieve a mean of 1000 photons per pixel in the white channel (sum of RGB channels) based on the training set statistics.
The mutual information between noiseless and noisy measurements was calculated using an analytic Gaussian approximation to ensure differentiability of the loss function with respect to the learnable parameters [53]. At each optimization step, we process a batch of 2304 measurement patches ($24\times 24$ pixels each), selected randomly from the Gehler-Shi dataset. These measurements were vectorized to compute their covariance matrix, whose log eigenvalues are used to estimate $H(\mathbf{Y})$. For numerical stability, any non-positive eigenvalues were set to $10^{-8}$. The batch size was chosen to be 2304, 4 times the number of pixels in a patch (576 pixels), as we found a batch of at least this size produces the most stable estimation. $H(\mathbf{Y}|\mathbf{X})$ is calculated using the analytic formula detailed earlier (Sec. 23).
For the IDEAL optimization, we designed an $8\times 8$ repeating color filter array (CFA) pattern with four possible filters at each pixel: red, green, blue, and white. To learn which filter to place at each location, we used a temperature-based softmax approach [53]:

### v5

@@ Acknowledgements
We conducted experiments using the Gehler-Shi dataset [gehlerBayesianColorConstancy2008a, shiReprocessedVersionGehler] (implied license via explicit permission to use), which comprises 568 high-quality natural images. The dataset was partitioned into 461 training, 51 validation, and 56 test images, with each image subdivided into $24\times 24$ pixel patches. To establish a consistent baseline for illumination, we computed a white channel by summing the red, green, and blue channels, then normalized the dataset to achieve a mean of $1000$ photons per pixel in the white channel, based on the training split mean. Signal-dependent Poisson noise was subsequently applied to all patches.
@@
For mutual information estimation, we sampled $100,000$ patches from the test set for fitting a PixelCNN, reserving $10,000$ for evaluation. Conditional entropy calculations were performed on clean images prior to synthetic Poisson noise application. We evaluated three distinct color filter array (CFA) designs: a Bayer pattern, a random configuration, and a learned pattern using the architecture in [chakrabartiLearningSensorMultiplexing]. For each CFA design, we selected the minimum information estimate across the 20 replicates (since our estimator provides an upper bound) and the best prediction performance (lowest negative log likelihood on test data) to characterize system performance. Each replicate took $\sim$3.5GB of memory. All replicates were trained in $\sim$5 hours on an Nvidia RTX A6000 GPU.
@@
Color image reconstruction was implemented using a bifurcated network architecture as proposed by Chakrabarti [chakrabartiLearningSensorMultiplexing]. This approach processes $24\times 24$ pixel measurement patches through two parallel paths. The first path performs multiplicative operations in log-space with linear mixing, while the second employs convolutional layers with $128$ filters. Each path generates $24$ estimates per color value, which are then combined through learned weights for final reconstruction.
@@
For our site selection experiments, we generated a dataset of 100,000 synthetic black holes based on the physical models for black hole dynamics from [levisInferenceBlackHole2021]. Each synthetic black hole had a static envelope, right ascension, and source declination matching that of M87. The dynamic portion forming the accretion disk surrounding each black hole was modeled as a Gaussian random field generated by a different solution to an underlying anisotropic spatio-temporal diffusion partial differential equation with fixed diffusion and advection fields. Each black hole was formed by applying a Gaussian random field as a multiplicative perturbation to the static envelope, and normalized to have total flux of 1 Jansky. To simulate black hole imaging we selected combinations of four telescopes from the eight original telescope locations used in the Event Horizon Telescope 2017 array. We also included comparisons between reconstruction quality and information estimates for all 70 possible combinations of four telescopes (Fig. S20). For each combination we simulated radio telescope measurements including thermal noise using the methods described in [chaelInterferometricImagingDirectly2018].
@@
Black hole images were reconstructed from simulated radio telescope measurements using a standard regularized maximum likelihood method with total variation regularization [chaelInterferometricImagingDirectly2018], with hyperparameters matching the implementation in [linImagingEvolvingBlack2024]. Images were reconstructed for a 2,000 measurement subset of the test set used for information estimation. Each reconstruction was performed on the CPU and took $\sim 10$s.
@@
The CIFAR10 dataset [krizhevskyLearningMultipleLayers] (implied license via permission to use) was used for the lensless imaging experiments. This dataset consists of 60,000 total images across 10 classes. To convert CIFAR10 image pixel values to synthetic photon counts, the dataset was scaled to have mean value equivalent to the desired photons per pixel. Encoded measurements for each lensless imaging system were generated by convolving (valid region, no zero padding) with the corresponding point spread function. The point spread functions for the one lens and four lens encoders were modeled as Gaussian-blurred points. For the diffuser encoder, the point spread function is a downsampled version of an experimentally captured point spread function from DiffuserCam [kabuliHighQualityLenslessImaging2023]. After convolution, synthetic Poisson noise was added to form the noisy encoded measurements.
@@
We analyzed single leukocyte images and corresponding protein expression measurements from the Berkeley Single Cell Computational Microscopy (BSCCM) dataset [pinkardBerkeleySingleCell2024] (CC0 1.0 Universal license). The dataset includes images acquired under multiple illumination conditions using an LED array microscope along with measurements of eight protein markers (CD3, CD19, CD56, CD123, HLA-DR, CD14, CD16, and CD45) obtained through antibody staining. To study performance at different signal levels, we simulated lower photon counts by adding synthetic Poisson noise to the original images.
@@
We developed a neural network to predict protein expression levels from single-cell images. The network uses a DenseNet121 [huangDenselyConnectedConvolutional2017] backbone pretrained on ImageNet, modified to accept single-channel input by averaging the first convolutional layer’s RGB weights. The backbone feeds into a global average pooling layer followed by eight parallel fully-connected networks, each with two hidden layers. These networks output parameters for Gaussian mixture models corresponding to each protein marker, enabling direct evaluation of the negative log likelihood for target protein levels. Training used the Adam optimizer with a learning rate of $5\times 10^{-5}$ and batches of 16 images. A composite loss function summed the negative log likelihood across protein markers, with missing values masked out. Training proceeded for 4,000 steps per epoch and employed early stopping when validation loss failed to improve for 20 epochs. Typical trainings took 1-2 days on NVIDIA TITAN Xp GPUs with 12GB memory. 2-3 networks were trained in parallel sharing a single GPU.
@@
S6.6 Information-Driven Encoder Analysis Learning (IDEAL)
Using the Gehler-Shi dataset [gehlerBayesianColorConstancy2008a, shiReprocessedVersionGehler] partitioned into training (461 images), validation (51 images), and test (56 images) sets, we extracted 100,000 patches of size $24\times 24$ pixels for training, along with 10,000 patches from the validation set. To establish consistent illumination conditions, the dataset was normalized to achieve a mean of 1000 photons per pixel in the white channel (sum of RGB channels) based on the training set statistics.
The mutual information between noiseless and noisy measurements was calculated using an analytic Gaussian approximation to ensure differentiability of the loss function with respect to the learnable parameters [chakrabartiLearningSensorMultiplexing]. At each optimization step, we process a batch of 2304 measurement patches ($24\times 24$ pixels each), selected randomly from the Gehler-Shi dataset. These measurements were vectorized to compute their covariance matrix, whose log eigenvalues are used to estimate $H(\mathbf{Y})$. For numerical stability, any non-positive eigenvalues were set to $10^{-8}$. The batch size was chosen to be 2304, 4 times the number of pixels in a patch (576 pixels), as we found a batch of at least this size produces the most stable estimation. $H(\mathbf{Y}|\mathbf{X})$ is calculated using the analytic formula detailed earlier (Sec. 24).
For the IDEAL optimization, we designed an $8\times 8$ repeating color filter array (CFA) pattern with four possible filters at each pixel: red, green, blue, and white. To learn which filter to place at each location, we used a temperature-based softmax approach [chakrabartiLearningSensorMultiplexing]: